Stormy Weather has no attractive investment opportunities. Its return on equity equals the discount rate, which is 10%. Its expected earnings this year are $4 per share. Complete the following table. (Do not round intermediate calculations. Enter the growth rate as a whole percent.):

Plowback Growth Rate Stock Price P/E Ratio Ratios 0 а. b. 0.40 0.80 C.
Answer #1

Following is the complete table

Plowback ratio ROE Growth rate EPS DPS Price P/E ratio
0 10.0% 0 x 10% = 0% 4 4 x (1-0) = 4 4 /(10%-0)= 40 40/4 =10
0.4 10.0% 0.4 x 10% = 4% 4 4 x (1-0.4) = 2.4 2.4/(10%-4%)= 40 40/4 = 10
0.8 10.0% 0.8 x 10% = 8% 4 4 x (1-0.8) = 0.8 0.8 /(10%-8%)= 40 40/4 = 10

Given information:

  • Plowback ratio
  • ROE
  • EPS

Calculations:

  • Growth rate = ROE x Plowback ratio
  • DPS= dividend per share
  • DPS = EPS x (1-Plowback ratio)
  • Price = DPS ROE - Growth rate
  • The formula for price is used above because we are given expected earnings and not current earnings so automatically the dividend calculated from that is expected dividend. Otherwise, had we been given the current earnings, we would have multiplied the DPS by (1+ growth rate)
  • P/E Ratio = Price/EPS